OpenStudy (anonymous):

find a closed form for the generating function of sequence: 1,0,-1,0,1,-1,0,1.... anyone explain me, please

OpenStudy (amistre64):

what definess a "closed form"

OpenStudy (anonymous):

I think it's just notation of form, ignore it, i just copy exactly what the problem is, in any easier problem, it is just a form

OpenStudy (anonymous):

or, it stops at 1/something, not expand as\[\sum_{k-0}^{\infty}something\]

OpenStudy (amistre64):

is this sequence the coeffs of a power series? or are we talking about a function that generates this sequence?

OpenStudy (amistre64):

for example: cos(x * pi/2); x=0,1,2,3,... would create the sequence 1,0,-1,0,1,....

OpenStudy (amistre64):

http://mathworld.wolfram.com/GeneratingFunction.html im trying to reconcile with this

OpenStudy (anonymous):

What came to mind for me was the real part of \(i^n\) for \(n=0,1,2,\ldots\), so\[\Re(i^n)=\frac{i^n+(-i)^n}{2} \text{ for } n=0,1,2,\ldots\]

OpenStudy (amistre64):

Definition: Given a sequence a0 , a1 , a2 , …, we define the generating function of the sequence {an} to be the power series G(x) = a0 + a1 x + a2 x 2 + ... .

OpenStudy (anonymous):

This reduces to amistre64's answer, just for the record.

OpenStudy (amistre64):

1,0,-1,0,1,-1,0,1.... \[1+0x-1x^2+0x^3+1x^4+0x^5+1x^6+....\] \[1-x^2+x^4-x^6+x^8-x^{10}+....\]

OpenStudy (amistre64):

yay!! :) my answer is a reduction

OpenStudy (amistre64):

\[\sum_{n=0}^{\infty}(-1)^n x^{2n}\]

OpenStudy (amistre64):

does that sound about right?

OpenStudy (anonymous):

It sounds right. but how to find it out? the problem comes from my yesterday test, i didn't have the solution yet. so that why i need the explanation

OpenStudy (amistre64):

denoting the summation is not closed form; closed form is putting this as a normal looking function

OpenStudy (anonymous):

but in my lecture the closed form should be the form of coefficient of x. like (-1)^n / (1 -x^2)

OpenStudy (anonymous):

and I must put everything in mathematical way, I mean show my work how can i get it form

OpenStudy (amistre64):

closed form doesnt have the "n" in it .... unless maybe its finite.

OpenStudy (anonymous):

yes, it is,

OpenStudy (amistre64):

the given is spose to represent 1/(x^2+1) when |x| < 1 which reminds me of finding the interval of convergence

OpenStudy (anonymous):

oh yeah. you are right. not n in it, n is in sum form. sorry, my bad

OpenStudy (amistre64):

lets take what we have so far: \[\sum(-1)^nx^{2n}\] and take the limit of the ratio as n approaches infinity \[\lim_{n\to~inf}\frac{(-1)^{n+1}x^{2(n+1)}}{1}*\frac{1}{(-1)^nx^{2n}}\] \[\lim_{n\to~inf}(-1)^{n-n+1}x^{2(n+1)-2n}\] \[\lim_{n\to~inf}-x^{2}\] pull out all the parts that dont have to deal with n \[|-x^2|~\lim_{n\to~inf} n^0=|x^2|\] i hope im recalling this correctly

OpenStudy (anonymous):

Thanks, I understand what you are doing. you put everything in converge an alternating function and find the result. but this part comes from discrete math, and your result from the previous reply is right. the solution for the problem is 1/1+x^2

OpenStudy (anonymous):

since i take long division of it , i get the coefficient of x is exactly the sequence i have from the problem.

OpenStudy (amistre64):

im glad you can make some sense of it :)

OpenStudy (anonymous):

\[\frac{ 1 }{1+x^2 }=\sum_{k=0}^{\infty}f(k)x^k\] and find out sum f(k)=1 +0+(-1)+0+1+0+(-1) and so on thanks a lot

OpenStudy (anonymous):

that makes sense to me, however, i screw it up. the good news is I understand now. thank thank thanks

OpenStudy (experimentx):

i think this does not converge, you can't write that sum = 1/2 based on this generating function.

OpenStudy (amistre64):

thats a relief :) it converges for |x| < 1

OpenStudy (amistre64):

for the interval (-1, 1) it is equivalent to 1/(1+x^2)

OpenStudy (anonymous):

I will have the solution tomorrow. if it is not as we do now, i will let you know, we can discuss more. now i have to go to school . again. thanks a lot. see you

OpenStudy (experimentx):

I think this is related to Cesaro summability, The Cesaro sum of alternating series is 1/2

OpenStudy (anonymous):

Don't forget about the geometric series. \[1-x^2+x^4-x^6+\cdots = 1+(-x^2)+(-x^2)^2+(-x^2)^3+\cdots = \frac{1}{1+x^2}\] for \(|-x^2|<1 \Rightarrow |x| < 1\).

OpenStudy (amistre64):

ahh, that makes the cobwebs a little clearer, thnx :)